OpenStudy (anonymous):

I'm quite confused. A right triangle has legs e and f and hypot. h. e = 10. The angle opposite leg f is 35*. What is the length of the hypotenuse?

OpenStudy (anonymous):

I know I need to use the pythagorean theorem when I get the side length, and I know how to get the side length. but do I use sine, cosine, or tangent?

OpenStudy (anonymous):

Tangent

OpenStudy (anonymous):

Can you explain, please?

OpenStudy (anonymous):

I'm confused when finding which one to use.

OpenStudy (anonymous):

Tangent is the only one that involves finding the hypotenuse, using both legs .

OpenStudy (anonymous):

Ah, ok. Can you wait around so you cam confirm if my answer is right?

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

What did you get ?

OpenStudy (anonymous):

Working it out right now.

OpenStudy (anonymous):

You can go somewhere else if you'd like, i can tag you.

OpenStudy (anonymous):

Tan35 = o/h

OpenStudy (anonymous):

Its fine .

OpenStudy (anonymous):



OpenStudy (anonymous):

Tan35 = o/a not o/h. Tan35 = f/10

OpenStudy (anonymous):

Wait the angle is 35 degrees ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

What is e then ?

OpenStudy (anonymous):

e = 10

OpenStudy (anonymous):

Tan35 * 10

OpenStudy (anonymous):

So e and f is 10 ?

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

F = unknown E = 10 H= unknown Angle = 35

OpenStudy (anonymous):

Oh i got it okay

OpenStudy (anonymous):

0.7002 * 10 = 7.002

OpenStudy (anonymous):

F = 7

OpenStudy (anonymous):

10^2 + 7^2 = c^2

OpenStudy (anonymous):

170 = c^2

OpenStudy (anonymous):

sqrt170

OpenStudy (anonymous):

Which cannot be reduced. Or equals 13.04

OpenStudy (anonymous):

So then H = 13.03

OpenStudy (anonymous):

Not 13.04

OpenStudy (anonymous):

How does it look?